OpenStudy (anonymous):

Suppose an equilateral triangle has sides given by 3y - 4 5y - 18 2y + 3 What is the length of each side of the triangle?

OpenStudy (anonymous):

The sides are all equal

OpenStudy (nurali):

3y - 4=5y - 18 14=2y y=7 so 3(7) - 4=17 5(7) - 18=17 2(7)+ 3=17 so the side lengths are 17

OpenStudy (whpalmer4):

Set the expressions for any two sides equal to each other and solve for y. Then use that value to find the length of each side.